OpenStudy (anonymous):

SAT question: If x=5y^(t+1) and y does not equal zero, then x/y= ?

OpenStudy (anonymous):

one not very satisfactory answer is \(\frac{x}{y}=5y^t\)

OpenStudy (anonymous):

the answer choices are: a. 5y b. y^t c. y^t/5 d. 5y^t e. 5y^(t+2) sorry, completely forgot about typing the answer choices @satellite73

OpenStudy (anonymous):

then i would go with D

OpenStudy (mertsj):

Just like @satellite73 said....D

OpenStudy (anonymous):

\[x=5y^{t+1}=5y^t\times y\] so \[\frac{x}{y}=5y^t\]

OpenStudy (anonymous):

@satellite73 where does the +1 go when you split it up?

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

\(b^{n}{b^m}=b^{m+n}\)

OpenStudy (anonymous):

so \(y^{t+1}=y^ty\)

OpenStudy (anonymous):

@satellite73 Why is it 5y^t and not 5y^t x y ?

OpenStudy (anonymous):

you lost me

OpenStudy (anonymous):

And how is that the answer overall? I don't see the link between the answer and x/y. Sorry for the hassle

OpenStudy (anonymous):

ok lets go slow

OpenStudy (anonymous):

\[x=5y^{t+1}\] is what is given

OpenStudy (anonymous):

that is the same as saying \[x=5y^t\times y\] since \(y^t\times y=y^{t+1}\)

OpenStudy (anonymous):

then if you divide both sides by \(y\) you get \[\frac{x}{y}=\frac{5y^t\times \cancel{y}}{\cancel{y}}=5y^t\]

OpenStudy (anonymous):

@satellite73 How do you know that 5y^t x y is the numerator?

OpenStudy (anonymous):

OH okay NEVERMINd

OpenStudy (anonymous):

ok